OpenStudy (anonymous):

solve for x: sec(x) * cos ((pi/2) - x) = -1

OpenStudy (anonymous):

s\[secx \cos(\frac{ \Pi }{ 2 }-x) = -1\rightarrow secx sinx = -1 \rightarrow \frac{ sinx }{ cosx }=-1\] tanx = -1 \[tanx = \tan 135^{0} \rightarrow x = 135^{0}\]